OpenStudy (anonymous):

Please Help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!(√12+2√8)(√12-2√8)

OpenStudy (anonymous):

yes we can do this

OpenStudy (anonymous):

it is like \[(a+b)(a-b)=a^2-b^2\] with \[a=\sqrt{12},b=2\sqrt{8}\]

OpenStudy (anonymous):

you get \[(\sqrt{12})^2-(2\sqrt{8})^2=12-4\times 8=12-32=-20\]\]

OpenStudy (anonymous):

That makes sense. Thanks a lot.